OpenStudy (vishweshshrimali5):

3,7,11.....407 nd 2,9,16....709 find num of common terms

OpenStudy (vishweshshrimali5):

answer it fast

OpenStudy (anonymous):

3,7,11.....407 difference is 4 There fore nth term is 4n + c [some constant] Constant can be found by looking at the 'zero' term First term is 3 therefore zero term is -1 [by ensuring the difference between them is 4] Nth term is 4n -1 Now make an equation 407 = 4n -1 408 = 4n 102 = n Therefore there are 102 terms in this sequence Try the other for your self and see how you get on. I will be around if you need help. Ciao Gianfranco

OpenStudy (vishweshshrimali5):

Now, for the second ap. 2,9,16,...................., 709 d = 7; a = 2; n= ? 709 = 2 + (n-1) 7 => n = 102

OpenStudy (vishweshshrimali5):

Now what to do ?

OpenStudy (anonymous):

both are arithmetic progressions. 2,9,16....709 \[first term,t_1=2 \space \space \space \space last term, t_n=709\] common difference= difference between any 2 consecutive terms.\[t_2-t_1=9-2=7\] \[no\space of \space terms\space \space ,n =\frac {(t_n-t_1)}{d} +1\] here, \[n=\frac{709-2}{7} +1=\frac{707}{7} +1=101+1=102\]

OpenStudy (vishweshshrimali5):

I have calculated the no. of terms in both case and it is 102

OpenStudy (vishweshshrimali5):

Now what to do I have to calculate the number of common terms in both ap.

OpenStudy (anonymous):

Hmm, the first set of terms are of form 7k +2 and the second set are 4k +3, is that right?

OpenStudy (anonymous):

When does 4k+3 = 7k +2? k= 1/3 not an integer, so no common terms.

OpenStudy (vishweshshrimali5):

Ok thanks for help estudier